A man expects to receive Php 125,000 in eight years. How much is that money worth now considering an interest rate of 12% compounded monthly?


1
Expert's answer
2022-01-09T13:41:27-0500

"A =P(1+\frac{r}{n})^{n×t}"

For compounded monthly n = 12

t = 8 years

r = 12%

A= Php 125,000


"125000 = P(1+\frac{0.12}{12})^{12×8}"


"125000=P(1.01)^{96}"


"125000=P(2.59927)"


"\frac{125000}{2.59927}=P\frac{2.59927}{2.59927}"


"P=Php 48090.43"

Money worth now is Php 48090.43